OpenStudy (anonymous):

For the two types of circuits (Series, and parallel) come up with rules in regards to current, voltage, and resistance when adding more bulbs

OpenStudy (blues):

Here is a place to start: The current through resistive elements in series is equal. Elements cannot accumulate current inside them - that is, all the current which flows in must flow out. That is true for both resistors in the circuit. The voltage across resistive elements in parallel is also equal. It works because if you were to replace the two resistors by modelling with a single resistor, the voltage wouldn't be changed. Is that helpful?

OpenStudy (anonymous):

im really confused... for resistance i have Series: As resistance increases, current decreases Parallel: As you add resistance to the circuit only the wire with the resistor connected to it has a decrease in current Current Series: The current in the circuit is equal to the volatge supplied by the entire resistance of the current Parallel: Current depends on the individual path and the............????? Voltage: Series: The total voltage across a series, circuit is divided among the individual electrical devices Parallel: ????????????????????????????

OpenStudy (blues):

I = V/R where I is current, V is voltage and R is resistance is a general relationship which holds for all resistive devises and anything which can be modelled with a resistive device. You have interpreted the formula correctly: when resistance increases, that is the denominator gets stronger, the current through the device decreases. Is that part clear?

OpenStudy (anonymous):

yes it is very clear

OpenStudy (ujjwal):

when connected in parallel, its like you are adding channels for current to flow.. and current will definately choose the path it finds easiest to flow in.. that is with least resistance.. so max current flows through least resistance..

OpenStudy (anonymous):

what do i say for the resistance in a parallel circuit?

OpenStudy (blues):

That equation is called the "V-I relationship" for the resistor. It holds no matter what kind of circuit (series, parallel) the resistors are in. The resistance, R, is just a property of the device - like a bulb has a certain resistance - and it is always constant. The I and the V depend on the circuit topology - what current sources and voltage sources are already in it...

OpenStudy (anonymous):

so would u just put R=V/I for the resistance in a parallel and series circuit?

OpenStudy (ujjwal):

Potential difference between each resistor in parallel is equal. now we have relation I=V/R.. sice V is equal for all resistors in parallel, I is inversely proportional to R.. so, least the resistance, max the current..

OpenStudy (anonymous):

thats what u would put?

OpenStudy (blues):

No, but you can use that relationship to derive the total resistance of two resistors which are in parallel and two which are in series. When resistors R1 and R2 are in parallel, the total resistance Rt = (R1 * R2) / (R1 + R2). When resistors R1 and R2 are in series, the total resistance is just additive. That is, Rt = R1 + R2.

OpenStudy (anonymous):

I am so confused, I dont know what to write :( I need to fill in the box on on my paper

OpenStudy (anonymous):

I realize what u are saying and i get it, i just dont know what to write

OpenStudy (blues):

I am glad that you get what I'm saying (that is the important part, really), but I don't know what sort of answer the box on your paper wants either. Does it want the formula, in which case you'd type in RT = R1 + R2. Or does it want words like, "The total resistance of two resistors in series is the sum of those resistors?"

OpenStudy (anonymous):

it wants words

OpenStudy (blues):

Then you should type out the equations in words. I just gave you the one sentence interpretation for resistors in series. What do you think the sentence would be for resistors in parallel? I apologize for getting series and parallel mixed up in my initial post.

OpenStudy (anonymous):

it said, come up with rules in regards to current, voltage, and resistance when adding more bulbs

OpenStudy (blues):

Well those two equations are your basis for those rules. For example, adding more bulbs (i.e., resistors or resistive elements) in series will increase the total resistance of the circuit. As the total resistance goes up, look at what effect it has on the I = V / R relation: current goes down. You can do so something similar for voltage by rearranging I = V / R so that it becomes an equation for V in terms of I and R. And you can apply the same sort of analysis to the equation for resistors in parallel.

OpenStudy (anonymous):

what about the current in a series vs parallel?

OpenStudy (blues):

Compare the two equations for the resistances Rt = R1 + R2 to Rt = (R1 * R2) / (R1 + R2). You can generalize those equations for n resistors of R resistance each: Rt = R * n and Rt = R^n / R*n for series and parallel resistors. You can compare the effect that increasing the number of resistors, n, has on the total resistance of each equation. One is linear, the other is exponential. So you can talk about what effects that difference in resistance increase has on current and voltage through the element.

OpenStudy (anonymous):

what about voltage?

OpenStudy (blues):

Well that's for you to figure out. Having general expressions for the total resistance of a series and parallel circuit with n resistors - the ones in the last response - is the key part. You can plug those expressions for Rt in terms of n into I = V / R and see what happens to I and V as n changes.

OpenStudy (anonymous):

the greater the resistance and current the greater the voltage for a series and parallel circuit

OpenStudy (blues):

What you need to do is take I = V / R. For this you need to substitute your two general equations for Rt (series and parallel) into the equation for R. These two equations are both resistance in terms of n - the R in them is just a constant. What you really get are expressions for I in terms of V, a variable and n. What happens to I as V is held constant and n increases and decreases? Then rearrange I = V / R so that it is an expression for V in terms of I and R. Do the same thing: plug your two general equations for R in terms of n into your equation for V. What happens as I is held constant and n changes? The first step will give you the effect on current. The second step will give you the effect on voltage.

OpenStudy (blues):

Any input, ujjwal?

OpenStudy (anonymous):

Im right though...

OpenStudy (blues):

To assess the effects of resistance on voltage and current, you need to do two separate steps. First you need to hold voltage constant and see what effect increasing and decreasing resistance has on current. Then you need to rearrange I = V / R so that it is an expression for voltage in terms of current and resistance. Then do the same thing for voltage that you did for current: namely, hold current constant and see what effect varying the resistance has on voltage. I have to go soon...

OpenStudy (anonymous):

the greater the resistance and current the greater the voltage for a series and parallel circuit V=I*R

OpenStudy (blues):

Yes, that is quite right.

OpenStudy (anonymous):

as you decrease resistance the current increases

OpenStudy (anonymous):

I=V/R

OpenStudy (blues):

Yes, exactly.

OpenStudy (blues):

All set?